An escalator has a slope of 3/4. After traveling forward 32 feet, the escalator is 24 feet above the floor. Write this word problem in point slope form .
Answer by TimothyLamb(4379)

for the point-slope form: y - y1 = m(x - x1)
we need a point: (x1, y1)
and the slope: m
---
the slope is given as 3/4
---
and a point is given as (32, 24).
but we can make things even easier by placing the foot of the escalator at (0, 0).
as you will see, it doesn't matter which point on the escalator we choose, the equations will be the same regardless.
---
the eqn with point (0, 0):
y - y1 = m(x - x1)
y - 0 = 3/4(x - 0)
---
y = 3/4(x)
---
the eqn with point (32, 24):
y - y1 = m(x - x1)
y - 24 = 3/4(x - 32)
y = 3/4(x) - 3/4(32) + 24
y = 3/4(x) - 24 + 24
---
y = 3/4(x)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php